You work for a company that has projected the following cash flows: Year 1: $2 million; Year 2: $3 million; Year 3: $4.5 million. Comparable companies have P/E ratios of 4x. What is the terminal value of the company? $4.5 million $9 million $9.5 million $18 million.
Answer #1

Terminal value of the company

Firstly, Calculate the Required Rate of Return (Ke)

Required Rate of Return (Ke) = 1 / P/E Ratio

= 1 / 4 Times

= 0.25 or 25%

Terminal value of the company

Terminal value = Cash flow in Year 3 / Required Rate of Return

= CF3 / Ke

= $4.50 Million / 0.25

= $18 Million

“Hence, the Terminal value of the company would be $18 Million”